TFY4155 Elektromagnetisme FY1003 Elektrisitet og magnetisme I Våren 2005

Curriculum (”Pencum”) in E. Lillestøl, O. Hunderi, J. R. Lien. Generell fysikk for universiteter og høgskoler. Bind 2 Varmelære og elektromagnetisme.

Universitetsforlaget 2001.

• 19: Electric charges and fields. The whole chapter.

• 20: Capacitors and capacitance. The whole chapter.

• 21: Current and resistance. 21.1 - 21.3.

• 22: DC-circuits. The whole chapter, but from ”Batterier” p 307 and down p 309 is not required.

• 23: Magnetic field. The whole chapter, except 23.8.

• 24: Faraday’w law. The whole chapter.

• 25: Inductance. The whole chapter.

• 26: Magnetic field in matter. 26.1 - 26.5.

• 27: Electric oscillations and AC current. Only 27.1

The curriculum in other books is as given in the final version of ”Framdriftsplan”. In addition, the general rule is that the curriculum is defined through the lectures and the exercises, including solutions.

Institutt for fysikk 04.05.2005 Jon Andreas Støvneng